A mutual fund charges a 2% front-end load. If an investor invest $2,000 in the fund and the net asset value per share (NAVPS) is $5,00, how many units can he buy?

a. 408

b.400

c.386

d.392
Answer #1

Option D is correct 392 UNITS

Step - 1 Calculation of Net Amount Invested after charging front end load

Investor Invested 2000 in the fund

Fornt end load = 2000 * 0.02 = 40

Net amount Invested for which investor will get the units = 2000 - 40 = 1960

Step 2 - Calculation of Units

No of Units = Net Amount Invested / NAVPS

No of Units = 1960 / 5 = 392

Therefore Investor can buy 392 Units